OpenStudy (anonymous):

1. Let f(x)=x^3-2x. Simplify the expression f(x+h)-f(x)/h

OpenStudy (anonymous):

Seems you are using definition of a derivative

Nnesha (nnesha):

first you should find f(x+h) so replace f(x) <---x in this equation by x+h

OpenStudy (anonymous):

You have to plug in (x+h) wherever there is an x in your function

OpenStudy (anonymous):

\[\frac{ (x+h)^3-2(x+h)-x^3-2x }{ h }\] so this

OpenStudy (anonymous):

You can do the simplification I think you just might have had trouble with the set up, I hope that helps.

Nnesha (nnesha):

\[\huge\rm f(\color{red}{x})= x^3-2x \] substitute x+h

Nnesha (nnesha):

\[\huge\rm \frac{ (x+h)^3-2(x+h)- \color{blue}{(}x^3-2x\color{blue}{)} }{ h }\] parentheses!!!

OpenStudy (anonymous):

Good call! :)

OpenStudy (anonymous):

so x^3-2x is the answer?? @Nnesha

Nnesha (nnesha):

nope how did you get that ?

OpenStudy (anonymous):

never mind, can you help me get the answer

Nnesha (nnesha):

okay first solve this (x+h)^3

OpenStudy (anonymous):

x+h^3

Nnesha (nnesha):

ahh nope (x+h)^3 is same as (x+h)(x+h)(x+h)

Nnesha (nnesha):

so either foil that or you can use this rule \[\rm (a+b)^3 = a^3 +3a^2b+3ab^2+b^3\]